# View-Centric Multi-Object Tracking with Homographic Matching in UAV Scene

### v2

@@ 3.4 Homography Matching Filter
3.5 Loss Design and Overall Tracking Algorithm
@@
where $\lambda$ is loss weight. We use cross-entropy loss both in $\mathcal{L}_{rec}$.
The proposed tracking algorithm can be summarized as Algorithm 1.
Input: An UAV vide sequence $\{I_{t}\in\mathbb{R}^{W\times H\times 3}\}^{T}_{t=1}$
Output: The tracked objects $T_{t}=\{B_{t},C_{t},ID_{t}\}$
   while $t<T$ do
     if $t$ is divisible by $h$ then
       Calculated $H_{t,t+h}$ of $I_{t},I_{t+h}$ directly by Homography Estimation.
     end if
     Derive $H_{t,t+1}$ from $H_{t,t+h}$ with Fast Homography Estimation algorithm.
     Input two adjacent frames $I_{t},I_{t+1}$.
     Finish object detection, obtain the detected boxes $\{B_{t},C_{t}\}$, $\{B_{t+1},C_{t+1}\}$.
     Calculate IoU association with Homography Matching Filter.
     Obtain ID embedding features $F^{ID}_{t},F^{ID}_{t+1}$, calculate the feature association.
     Update the ID feature $F^{ID}_{t+1}$ with View-Centric ID Learning.
     Ensemble IoU and feature association, perform Matching Algorithm.
   end while
Algorithm 1 HomView-MOT algorithm
4 Experiments
4.1 Datasets
VisDrone2019: The VisDrone2019 dataset [56] is applied for tracking and detection from UAV perspectives. For the MOT task, the VisDrone2019 dataset comprises a training set (56 sequences), a validation set (7 sequences), and a test set (33 sequences, with test-challenge containing 16 sequences and test-dev comprising 17 sequences). In each frame, every object is labeled with a bounding box, category, and tracking ID. The VisDrone2019 dataset encompasses ten categories: pedestrian, person, car, van, bus, truck, motor, bicycle, awning-tricycle, and tricycle. For the multi-object tracking evaluation, only five object categories are considered, namely car, bus, truck, pedestrian, and van.
@@ 4.1 Datasets
MOTDT [5] - -0.8 21.6 87 1196 44548 185453 1437 3609
SORT [2] 23.5 14.0 38.0 506 545 80845 112954 3629 4838
IoUT [3] 27.3 28.1 38.9 467 670 36158 126549 2393 3829
GOG [11] 2.0 28.7 36.4 346 836 17706 144657 1387 2237
MOTR [40] 7.5 22.8 41.4 272 825 28407 147937 959 3980
TrackFormer [24] 7.4 25 30.5 385 770 25856 141526 4840 4855
UAVMOT [22] 12.0 36.1 51.0 520 574 27983 115925 2775 7396
ByteTrack [42] 11.4 52.3 68.3 769 325 24156 110342 2230 4001
U2MOT [21] 19.4 52.3 69.0 - - - - 1052 -
HomView-MOT 20.8 54.2 75.1 870 224 31461 72600 1073 2832
Table 1: Comparison with state-of-the-arts on VisDrone test-dev dataset.
@@
CEM [25] -6.8 10.1 94 1062 64373 298090 1530 2835
SMOT [7] 33.9 45.0 524 367 57112 166582 1752 9577
GOG [11] 35.7 0.3 627 374 62929 153336 3104 5130
IOUT [3] 36.6 23.7 534 357 42245 163881 9938 10463
CMOT [1] 36.9 57.5 664 351 69109 144760 $1$ 3656
SORT [2] 39 43.7 484 400 33037 172628 2350 5787
DSORT [34] 40.7 58.2 595 338 44868 155290 2061 6432
MDP [36] 43.0 61.5 647 324 46151 147735 541 4299
UAVMOT [22] 46.4 67.3 624 221 66352 115940 456 5590
HomView-MOT 48.1 70.7 694 234 58834 117691 427 6003
Table 2: Comparison with state-of-the-arts on UAVDT test dataset.
UAVDT: The UAVDT dataset [8] is specifically utilized for vehicle object detection and tracking, featuring three categories: car, truck, and bus. For the MOT task, it is divided into a training set (30 sequences) and a test set (20 sequences). It exclusively focuses on a single category, car. The video frames have a resolution of 1080 × 540 pixels and encompass various common scenes, such as squares, arterial streets, and toll stations.
4.2 Implementation Details
To ensure a fair comparison, we follow the experimental setup of existing works such as UAVMOT [22]. Regarding training details, the total number of epochs is set to 30, with the learning rate decaying by a factor of 10 at 10 epochs and 20 epochs, respectively. The initial learning rate is set to 7e-5. For data augmentation, we utilize random cropping and random scaling (between 0.6 to 1.3). The experiments are conducted on two GeForce RTX GPUs with a batch size of 4. In the multiple loss functions, L1 loss is used to supervise the object’s width and height. Cross-entropy loss and triplet loss are employed for handling the object ID. Additionally, we utilize the standard cross-entropy loss to supervise the object heatmaps and view slot reconstruction. The loss weight $\lambda$ is set to 0.4.
@@ 4.2 Implementation Details
Baseline 2079 40.6
Baseline+IDFU[22] 937 43.8
Baseline+VCIL 790 51.9
Table 3: The ablation study of View-Centric ID Learning (VCIL) on VisDrone validation set.
@@
Baseline 2079 40.6
Baseline+AMF[22] 958 44.1
Baseline+HMF 816 53.7
Table 4: The ablation study of Homographic Matching Filter (HMF) on VisDrone validation set.
Figure 8: The robustness analysis of Fast Homography Estimation (FHE). Horizontal axis: Sampling interval $h$ in the FHE algorithm. Vertical axis: Tracking accuracy (IDF1) and speed (FPS) of HomView-MOT under different settings of $h$. When $h=1$, FHE degenerates into the ordinary HE algorithm, yielding high accuracy but at a slow speed. As $h$ increases, the tracker maintains relatively high accuracy while the speed significantly improves, confirming the value of the FHE approach. However, if $h$ is too large, the calculation of the Homography matrix becomes imprecise, which affects performance.
@@
4.3 Comparison with State-of-the-arts
We evaluate from the two most critical dimensions: accuracy and speed, on the VisDrone and UAVDT datasets.
VisDrone Following previous works[22, 21], we benchmark our method against previous approaches on the VisDrone2019 dataset for the MOT task. We train our model using both the training and validation sets and assess our approach on the VisDrone2019 test-dev set utilizing the official VisDrone MOT toolkit. As depicted in Table 1, our method attains a 54.2% score on MOTA and 75.1% on IDF1, surpassing existing techniques on the VisDrone2019 test-dev set.
UAVDT In addition to our evaluation on the VisDrone2019 dataset, we also benchmark our method against other approaches on the UAVDT test set for the MOT task. We train the proposed network on the UAVDT training set and evaluate our approach on the UAVDT test set. We present a range of metrics, including MOTA and IDF1, to compare the performance of our method with that of others. As shown in Table 2, our method registers a 48.1% score on MOTA and 70.7% on IDF1, achieving significantly enhanced results compared to existing methods.
4.4 Ablation Study
4.4.1 Effectiveness of View-Centric ID Learning
View-Centric ID Learning (VCIL) enables ID embedding to learn multi-view information, effectively enhancing the reliability of ID feature association and thereby bolstering the robustness of MOT algorithms in moving UAV scenarios. To verify the effectiveness of VCIL, we study its benefits on ID feature association using two metrics: IDs, IDF1. For fair comparison with previous methods, we follow the same baseline with [22]. Specifically, we evaluate the performance of the baseline, the baseline with the IDFU module introduced in [22], and the baseline with VCIL. The experimental results, as shown in Table 4.2, indicate that the inclusion of VCIL leads to a significant improvement across all metrics compared to both the baseline and IDFU. These results demonstrate that VCIL has a substantial positive impact on the ID feature association process, allowing the model to quickly acquire multi-view information of targets during tracking.
4.4.2 Effectiveness of Homography Matching Filter
To address the challenges posed by the irregular motion of targets in moving UAV scenarios, we present the Homography Matching Filter (HMF). To validate its effectiveness, we compare its performance with that of ordinary IoU association, the adaptive motion filter (AMF) proposed in [22], as well as against HMF itself. For fair comparison with previous methods, we follow the same baseline with [22]. Particularly, our HMF is a lightweight yet effective algorithm. The experimental results, as presented in Table 4.2, are based on two metrics: IDs, IDF1. It is evident that HMF demonstrates improved accuracy over both the baseline and AMF. Notably, in contrast to the manually intensive and complex "UAV motion mode selective + local relation filter" used by AMF, HMF is more lightweight and, while maintaining higher accuracy, also offers faster speed.
4.4.3 Robustness Analysis of Fast Homography Estimation
Within the design of HomView-MOT, the calculation of inter-frame Scene Projection through Fast Homography Estimation (FHE) is fundamental and critical. In our design, to ensure the overall efficiency of tracking, we compute the inter-frame Homography Matrix rapidly by using interval sampling coupled with FHE, based on the assumption that the scene can be reasonably considered to be moving linearly within a small sampling interval. This design ensures the efficiency and precision of our HomView-MOT. Generally, a larger sampling interval tends to increase tracking efficiency, but an overly large interval may weaken the assumption of "approximate linear motion within the interval," thus compromising tracking accuracy. Figure 8 displays the performance of tracking under different settings of $h$. It is observed that a smaller $h$ yields the highest accuracy but also reduces speed. As $h$ increases, we maintain a high level of accuracy while also achieving a higher speed, but a further increase in $h$ will lead to a noticeable decrease in accuracy. These experimental findings strongly support the aforementioned arguments.
4.4.4 Case Study
To better illustrate the effectiveness of the Homography Matching Filter, we visually present in Figure 9 the performance of HMF when the UAV-captured scene moves, causing the target trajectories to become irregular. Figure 9 depict three different UAV status changes: hovering, turning up, and turning left, showing comparison between two consecutive frames. It is evident that the objects have not moved significantly in the real world (as can be seen with reference to ground features such as lane lines), and when directly mapped, the IoU between the detection boxes of consecutive frames is very low, leading to potential match losses. However, with the mapping based on Homography, the detection boxes from frame to frame are closer together, yielding a higher IoU and thus achieving a more reliable IoU association.
5 Conclusion
Our work presents a pioneering HomView-MOT framework that successfully addresses the inherent challenges of multi-object tracking in moving UAV scenarios. By innovatively employing the principles of Homography and View-Centric learning, our approach effectively navigates the complexities introduced by UAV flight dynamics, such as irregular motion patterns and varying view perspectives. The Fast Homography Estimation (FHE) algorithm computes homography matrices efficiently, while our Homographic Matching Filter (HMF) and View-Centric ID Learning (VCIL) techniques facilitate accurate and robust tracking across changing scenes. The effectiveness of our framework is underscored by its state-of-the-art performance on the VisDrone and UAVDT datasets.
Limitation and Outlook. In fact, HomView-MOT is a universal framework designed for "multi-object tracking in moving scenarios" and is not limited to UAVs alone; in theory, it can be applied to a variety of industrial applications including different types of aircraft, handheld recording devices, underwater diving cameras, and more. Due to the current limitations of available open-source datasets, we have only been able to validate our framework in moving UAV scenarios. In the future, we look forward to gathering more datasets relevant to diverse scenarios in order to more broadly ascertain the efficacy and versatility of our method.

### v3

@@ III-F Homography Matching Filter
Algorithm 1 HomView-MOT Algorithm
1:An UAV video sequence $\{I_{t}\in\mathbb{R}^{W\times H\times 3}\}_{t=1}^{T}$
2:The tracked objects $T_{t}=\{B_{t},C_{t},ID_{t}\}$
3:while $t<T$ do
4:  if $t$ is divisible by $h$ then
5:    Calculate $H_{t,t+h}$ of $I_{t}$, $I_{t+h}$ directly by Homography Estimation.
6:  end if
7:  Derive $H_{t,t+1}$ from $H_{t,t+h}$ with Fast Homography Estimation algorithm.
8:  Input two adjacent frames $I_{t}$, $I_{t+1}$.
9:  Finish object detection, obtain the detected boxes $\{B_{t},C_{t}\}$, $\{B_{t+1},C_{t+1}\}$.
10:  Calculate IoU association with Homography Matching Filter.
11:  Obtain ID embedding features $F^{ID}_{t}$, $F^{ID}_{t+1}$, calculate the feature association.
12:  Update the ID feature $F^{ID}_{t+1}$ with View-Centric ID Learning.
13:  Ensemble IoU and feature association, perform Matching Algorithm.
14:end while
III-G Loss Design and Overall Tracking Algorithm
@@
where $\lambda$ is loss weight. We use cross-entropy loss both in $\mathcal{L}_{rec}$. The proposed tracking algorithm can be summarized as Algorithm 1.
IV Experiments
IV-A Datasets
IV-A1 VisDrone2019
: The VisDrone2019 dataset [9] is applied for tracking and detection from UAV perspectives. For the MOT task, the VisDrone2019 dataset comprises a training set (56 sequences), a validation set (7 sequences), and a test set (33 sequences, with test-challenge containing 16 sequences and test-dev comprising 17 sequences). In each frame, every object is labeled with a bounding box, category, and tracking ID. The VisDrone2019 dataset encompasses ten categories: pedestrian, person, car, van, bus, truck, motor, bicycle, awning-tricycle, and tricycle. For the multi-object tracking evaluation, only five object categories are considered, namely car, bus, truck, pedestrian, and van.
IV-A2 UAVDT
: The UAVDT dataset [10] is specifically utilized for vehicle object detection and tracking, featuring three categories: car, truck, and bus. For the MOT task, it is divided into a training set (30 sequences) and a test set (20 sequences). It exclusively focuses on a single category, car. The video frames have a resolution of 1080 × 540 pixels and encompass various common scenes, such as squares, arterial streets, and toll stations.
IV-B Implementation Details
To ensure a fair comparison, we follow the experimental setup of existing works such as UAVMOT [7]. Regarding training details, the total number of epochs is set to 30, with the learning rate decaying by a factor of 10 at 10 epochs and 20 epochs, respectively. The initial learning rate is set to 7e-5. For data augmentation, we utilize random cropping and random scaling (between 0.6 to 1.3). The experiments are conducted on two GeForce RTX GPUs with a batch size of 4. In the multiple loss functions, L1 loss is used to supervise the object’s width and height. Cross-entropy loss and triplet loss are employed for handling the object ID. For all main experiments on the VisDrone and UAVDT datasets, the value of $h$ was uniformly set to 10. Additionally, we utilize the standard cross-entropy loss to supervise the object heatmaps and view slot reconstruction. The loss weight $\lambda$ is set to 0.4.
TABLE I: Comparison with state-of-the-arts on VisDrone test-dev dataset.
@@ IV-B Implementation Details
MOTDT [56] ICME2018 - -0.8 68.5 21.6 87 1196 44548 185453 1437 3609
SORT [3] ICIP 2016 23.5 14.0 73.2 38.0 506 545 80845 112954 3629 4838
IoUT [57] AVSS2017 27.3 28.1 74.4 38.9 467 670 36158 126549 2393 3829
MOTR [5] ECCV2022 7.5 22.8 72.8 41.4 272 825 28407 147937 959 3980
TrackFormer [21] CVPR2022 7.4 25.0 73.9 30.5 385 770 25856 141526 4840 4855
UAVMOT [7] CVPR2022 12.0 36.1 74.2 51.0 520 574 27983 115925 2775 7396
ByteTrack [19] ECCV2022 11.4 52.3 - 68.3 769 325 24156 110342 2230 4001
OC-SORT [12] CVPR2023 - 39.6 73.3 50.4 - - 14631 123513 986 -
FOLT [58] MM2023 - 42.1 77.6 56.9 - - 24105 107630 800 -
GLOA [59] J-STARS2023 - 39.1 76.1 581 824 18715 158043 4426 -
U2MOT [4] ICCV2023 19.4 52.3 - 69.0 - - - - 1052 -
STDFormer [60] TCSVT2023 - 45.9 77.9 57.1 684 538 21288 101506 1440 -
DroneMOT [61] ICRA2024 - 43.7 71.4 58.6 689 397 41998 86177 1112 -
AHOR-ReID [62] TCSVT2024 28.1 42.5 - 56.4 - - 21447 109762 810 -
DFA-MOT [63] TCSVT2025 19.8 42.9 - 58.3 518 523 10760 119571 792 -
HomView-MOT - 20.8 54.2 75.7 75.1 870 224 31461 72600 1073 2832
IV-C Comparison with State-of-the-arts
Following previous works[7, 4], we benchmark our method against previous approaches on both the VisDrone2019 and UAVDT datasets for the MOT task. For VisDrone2019, we train our model using both the training and validation sets and assess our approach on the test-dev set utilizing the official VisDrone MOT toolkit. For UAVDT, we train the proposed network on the training set and evaluate on the test set. As depicted in Table I and Table II, our method achieves significantly enhanced results compared to existing methods.
IV-D Ablation Study
IV-D1 Ablation Study of Each Component
We delve into each proposed component of HomView-MOT the show the results in Table III. Specifically, we evaluate HomView-MOT with different settings, including “w/o FHE", “w/o VCIL" and “w/o HMF". Noted that, for the setting of “w/o FHE", we set the homography matrix of any two frames to be the homography matrix of adjacent key frames. As seen that all the components play an important role for the proposed tracking framework.
TABLE II: Comparison with state-of-the-arts on UAVDT test dataset.
@@ IV-D1 Ablation Study of Each Component
IOUT [57] AVSS2017 36.6 72.1 23.7 534 357 42245 163881 9938 10463
SORT [3] ICIP2016 39.0 74.3 43.7 484 400 33037 172628 2350 5787
DSORT [11] ICIP2017 40.7 73.2 58.2 595 338 44868 155290 2061 6432
CenterTrack [64] ECCV2020 40.9 - 59.9 624 221 41681 159478 434 2093
ByteTrack [19] ECCV2020 41.6 - 59.1 - - 28819 189197 296 -
FOLT [58] MM2023 41.6 - 48.5 - 68.3 36429 155696 338 -
OCSORT [12] CVPR2023 47.5 - 64.9 - 68.3 47681 148378 288 -
UAVMOT [7] ICCV 2023 46.4 72.7 67.3 624 221 66352 115940 456 5590
DistMOT [65] ICSIDP2024 40.4 - 57.2 - - - - 593 -
DFA-MOT [63] TCSVT2025 48.0 - 69.3 684 230 59218 119248 396 1269
HomView-MOT - 48.1 74.3 70.7 694 234 58834 117691 427 6003
TABLE III: Ablation study of each component on VisDrone.
FHE VCIL HMF MOTA $\uparrow$ (%) IDF1 $\uparrow$ (%) IDs $\downarrow$
$\checkmark$ $\checkmark$ $\checkmark$ 54.2 75.1 1073
$\checkmark$ $\checkmark$ 42.2 64.7 2350
$\checkmark$ $\checkmark$ 48.8 68.5 1566
$\checkmark$ $\checkmark$ 50.0 70.9 1322
IV-D2 Ablation Study on VCIL
View-Centric ID Learning (VCIL) enables ID embedding to learn multi-view information, effectively enhancing the reliability of ID feature association and thereby bolstering the robustness of MOT algorithms in moving UAV scenarios. To verify the effectiveness of VCIL, we study its benefits on ID feature association using two metrics: IDs, IDF1. For fair comparison with previous methods, we follow the same baseline with [7]. Specifically, we evaluate the performance of the baseline, the baseline with the IDFU module introduced in [7], and the baseline with VCIL. The experimental results, as shown in Table IV, indicate that the inclusion of VCIL leads to a significant improvement across all metrics compared to both the baseline and IDFU.
IV-D3 Comparison between VCIL and ReID methods
To quantitatively validate VCIL’s advantages, we compare it against two standard ReID-based MOT baselines on the VisDrone dataset: 1) DeepSORT-style ReID, uses a pre-trained ReID model for feature extraction; 2) FairMOT-style Joint Detection and ReID, integrates ReID learning into the detection head. The results shown in Table. V demonstrate VCIL’s superiority in handling view changes specific to UAV scenarios. VCIL significantly reduces identity switches (IDs) and improves IDF1, highlighting its ability to maintain ID consistency under challenging viewpoint variations where standard ReID methods fail.
Figure 8: Ablation Study on $h$ in FHE on VisDrone.
IV-D4 Ablation Study on $h$ in FHE
In order to ensure the overall efficiency of tracking, we compute the inter-frame homography Matrix rapidly by using interval sampling coupled with FHE, based on the assumption that the scene can be reasonably considered to be moving linearly within a small sampling interval. This design ensures the efficiency and precision of our HomView-MOT. Generally, a larger sampling interval tends to increase tracking efficiency, but an overly large interval may weaken the assumption of "approximate linear motion within the interval," thus compromising tracking accuracy. Figure 8 displays the performance of tracking under different settings of $h$. It is observed that a smaller $h$ yields the highest accuracy but also reduces speed. As $h$ increases, we maintain a high level of accuracy while also achieving a higher speed, but a further increase in $h$ will lead to a noticeable decrease in accuracy. These experimental findings strongly support the aforementioned arguments. Specifically, When $h=1$, FHE degenerates into the ordinary HE algorithm, yielding high accuracy but at a slow speed. As $h$ increases, the tracker maintains relatively high accuracy while the speed significantly improves, confirming the value of the FHE approach. However, if $h$ is too large, the calculation of the homography matrix becomes imprecise, which affects performance. Therefore, in all the experiments, we set $h=10$ to balance the accuracy and efficiency.
TABLE IV: Ablation study on VCIL on VisDrone.
@@ IV-D4 Ablation Study on $h$ in FHE
Same Baseline as [7] 2079 40.6
+IDFU[7] 937 43.8
+VCIL 790 51.9
IV-D5 Comparison between HMF and the existing motion filters
Motion-aware filters such as Kalman excel at modeling object dynamics because the camera’s perspective is stable. However, in moving-UAV scenarios, the camera’s ego-motion (e.g., abrupt turns, altitude changes) often dominates the observed motion of objects. This makes it difficult to disentangle an object’s true motion from the camera’s movement using traditional motion models alone. The proposed HMF is to first eliminate the dominant camera motion via homography projection, thereby recovering the object’s residual motion in a common reference plane. As shown in Table VI, the Kalman filter struggles with erratic UAV ego-motion, leading to frequent trajectory mismatches (high IDs). In contrast, HMF’s homography-based stabilization proves more robust to camera dynamics, achieving higher accuracy and fewer identity switches. This validates that in moving-UAV settings, compensating for camera motion geometrically (via HMF) is more effective than predicting object motion under unstable camera dynamics.
TABLE V: Comparison between VCIL and ReID methods on VisDrone.
@@ IV-D5 Comparison between HMF and the existing motion filters
w/ DeepSORT ReID [38] 2911 38.3
w/ FairMOT-style ReID [33] 1033 41.7
w/ VCIL 790 51.9
TABLE VI: The Closer Observation on HMF on VisDrone.
[…]
Same Baseline as [7] 2079 40.6
+AMF[7] 958 44.1
+HMF 816 53.7
IV-D6 Case Study
To better illustrate the effectiveness of the HMF, we visually present in Figure 9 the performance of HMF when the UAV-captured scene moves, causing the target trajectories to become irregular. Figure 9 depict three different UAV status changes: hovering, turning up, and turning left, showing comparison between two consecutive frames. It is evident that the objects have not moved significantly in the real world (as can be seen with reference to ground features such as lane lines), and when directly mapped, the IoU between the detection boxes of consecutive frames is very low, leading to potential match losses. However, with the mapping based on homography, the detection boxes from frame to frame are closer together, yielding a higher IoU and thus achieving a more reliable IoU association.
IV-D7 Discussion of the generalization capability
The HomView-MOT framework is fundamentally designed to handle challenges that are universal across moving-camera scenarios, which naturally provides a degree of robustness against domain shifts. The HMF does not rely on learning dataset-specific motion patterns. Instead, it uses scene homography, a geometric property, to project object bounding boxes into a common view for association. Therefore, even if the UAV’s flight pattern (e.g., more aggressive turns) or the environment (e.g., from an urban grid to a highway) changes significantly, the underlying geometric principle remains valid. This makes HMF less sensitive to domain shifts in motion dynamics compared to Kalman-filter-based approaches that learn from training data.
V Conclusion
We propose the HomView-MOT framework that successfully addresses the inherent challenges of multi-object tracking in moving UAV scenarios. By employing the principles of homography and view-centric learning, our approach effectively navigates the complexities introduced by UAV flight dynamics. The FHE algorithm computes homography matrices efficiently, while our HMF and VCIL techniques facilitate accurate and robust tracking across changing scenes. The effectiveness of our framework is underscored by its state-of-the-art performance on the VisDrone and UAVDT datasets.
Limitation. The core design of HomView-MOT is not exclusively limited to UAV scenarios. The VCIL and HMF modules are designed to address the fundamental issue of scene motion, which is a common challenge in all moving-camera tracking. However, performance gains may vary in non-UAV contexts. In scenarios with milder and more predictable homography changes, the performance improvement offered by the HMF might be less dramatic than that in highly dynamic UAV flights, as the standard IoU association may already be reasonably effective. Nevertheless, the VCIL module’s ability to robustly handle view changes could still contribute to the tracking robustness.